# Ambient Noise Full Waveform Inversion with Neural Operators

### v3

@@ 1 Introduction
Seismic tomography for sedimentary basins is important for assessing earthquake hazards considering the fact that the basins can trap and amplify seismic waves, leading to stronger and longer ground shaking. Fine-scale subsurface structures are typically determined by active source surveys, which often operate at high frequencies but are expensive and limited by environmental impact issues. Alternatively, seismic data from passive sources, such as ambient noise and earthquakes, can be more readily acquired and can also compensate for the absence of low-frequency information in active source data. Many have shown in seismic interferometry that the cross-correlation between diffuse waveforms recorded at two stations approximates the elastodynamic Green’s function (Wapenaar, 2004; Wapenaar et al., 2010). The so-called empirical Green’s function (EGF) created from ambient noise is more commonly used by ray theory-based tomography methods (Guo et al., 2015; Lin et al., 2008; Yao et al., 2010; Zheng et al., 2011), which do not account for the finite-frequency effects in heterogeneous media. Full waveform inversion (FWI), which accounts for the full physics of wave propagation, can reveal more accurate and detailed subsurface structures (Liu et al., 2017, 2023a; Maguire et al., 2022; Sager et al., 2020; Wang et al., 2021a; Zhang et al., 2021).
@@
Despite its significant advantages, FWI can be time-consuming and memory-intensive for large problems, which hinders its broader applicability. Moreover, the adjoint-state method (Fichtner et al., 2006a; Liu & Tromp, 2006; Plessix, 2006; Tromp et al., 2005), which is conventionally used for FWI, requires the initial model to be sufficiently close to the true Earth to avoid being trapped in local minima (Gauthier et al., 1986; Fichtner et al., 2006b; Panning et al., 2009; Virieux & Operto, 2009). In ambient noise tomography (ANT), the adjoint-state method typically relies on a model derived from surface wave dispersion analysis and ray-based approximations as a starting point (Chen et al., 2014; Liu et al., 2023a; Maguire et al., 2022; Zhang et al., 2018, 2020). To tackle these challenges, recent studies have shown that neural operators (Li et al., 2020a, b; Azizzadenesheli et al., 2024) can solve the wave equations orders of magnitude faster than traditional numerical solvers (Yang et al., 2021, 2023; Zou et al., 2024). In deep learning, an analogue to the adjoint-state method is known as the reverse-mode automatic differentiation (AD) (Baydin et al., 2018; Rumelhart et al., 1986). While these two methods have been shown to be mathematically equivalent (Richardson, 2018; Wang et al., 2021b; Zhu et al., 2021), AD, which underpins the modern deep learning platform PyTorch (Paszke et al., 2019), remains aligned with advances in state-of-the-art optimization techniques. For example, methods such as mini-batching (LeCun et al., 2002), stochastic gradient descent (SGD) (Bottou, 2010), and adaptive moment estimation (Adam) (Kingma & Ba, 2014) have been developed to better navigate non-convex loss landscapes and escape poor local minima. These methods have prevailed for over a decade in the machine learning community, while remaining relatively underexplored in the classic FWI world. Only in recent years have they begun to attract attention (Bernal-Romero & Iturrarán-Viveros, 2021; Mao et al., 2025; Richardson, 2018; Sun et al., 2020; Thrastarson et al., 2022; van Herwaarden et al., 2020). Therefore, in alignment with the rapid advancements in optimization techniques, neural operators have the potential to alleviate — though not fully resolve — the cycle-skipping problem in FWI, making it possible to start from a simpler velocity model (e.g., a 1D model).
@@
The primary advantage of neural operators in solving partial differential equations (PDEs) over other prevailing machine learning approaches, such as physics-informed neural networks (PINNs) (Raissi et al., 2019), is that they are trained to learn the solution operator for an entire family of PDEs instead of a specific instance. This means that a neural operator has the potential to generalize to arbitrary PDE coefficients (e.g., velocity structures) if the training instances are ideally sampled — for example, from random fields (Yang et al., 2021, 2023; Zou et al., 2024). The training is a one-time effort for the forward process and there is no further training for the neural operator in the inversion stage, where the velocity parameters are updated with gradients computed from AD. This is similar to traditional FWI but replaces the adjoint simulation with back-propagation (LeCun et al., 1988).
@@
In recent years, neural operators have been increasingly applied to seismic wave propagation modeling. They have been used for modeling acoustic (Yang et al., 2021), elastic (Li et al., 2025; Yang et al., 2023), and viscoelastic (Wei & Fu, 2022) waves. Most applications are limited to 2D due to computational and memory constraints. However, Lehmann et al. (2024, 2025) circumvented the 3D computational bottleneck by predicting only the surface ground motion, while Zou et al. (2024) modeled the full 3D wavefield by parameterizing the operator in the frequency domain, thereby eliminating the need to model the time dimension. Learning frequency-domain solutions offers memory efficiency for neural operators and has attracted growing interest (Cheng et al., 2025; Huang & Alkhalifah, 2025; Kong et al., 2025; Li et al., 2023a; Zhang et al., 2023). Among these, Huang & Alkhalifah (2025) and Cheng et al. (2025) proposed learning the residual (scattered) wavefield instead of the full wavefield to address the point-source singularity. Current research has primarily focused on synthetic data, while the application of neural operators to real seismic data remains largely unexplored.
@@
Our contributions are as follows. We demonstrate the first application of neural operators for ambient noise tomography on a real seismic dataset, specifically the Basin Amplification Seismic Investigation (BASIN) survey (BASIN, 2018; Clayton et al., 2019), which deployed linear nodal arrays in the northern Los Angeles (LA) basins. We provide a trained neural operator that is applicable without retraining to any linear array in the study area, or potentially other problems with a comparable spatial footprint. With the trained neural operator, FWI can be performed two orders of magnitude faster than the conventional adjoint method. Our tomography results are consistent with previous studies in the same region (Ghose et al., 2023; Li et al., 2023b; Villa et al., 2023; Zou & Clayton, 2024) without the strong dependence on priors. The proposed method can be scaled to 3D, provided that sufficient computing resources are available.
@@ 2.1 Helmholtz Neural Operator with Automatic Differentiation
Based on the universal approximation theorem, neural operators can approximate arbitrary nonlinear continuous operators (Hornik et al., 1989; Kovachki et al., 2023). The solution operator can be parameterized much more efficiently in memory when solving Equation 2 in the frequency domain (i.e., the Helmholtz equation) than in the time domain (Huang & Alkhalifah, 2025; Kong et al., 2025; Zou et al., 2024). This is because different frequency components can be solved individually and, thus, parallelized. In this study, we employ the Helmholtz Neural Operator (HNO) proposed by Zou et al. (2024). The HNO computes the solution for each frequency component independently by defining the batch of data at the frequency level. In other words, a batch can contain multiple frequency components due to data shuffling. We use a single HNO to learn the entire frequency band of interest (0.10.10.1 to 0.50.50.5 Hz), although Kong et al. (2025) found that using multiple models to handle different frequencies slightly improves performance. Solutions in the time and frequency domains can be related via the (inverse) Fourier transform. We construct a U-shaped HNO comprising $L=8$ inner layers of Fourier Neural Operators (FNOs) (Li et al., 2020a), followed by a Graph Neural Operator (GNO) (Li et al., 2020b), to predict the waveforms at the free surface (data, denoted by $\mathbf{\hat{d}}$):
@@
where $\mathbf{a}$ is the input including $V_{P}$, $V_{S}$, the source location, and a constant function indicating the frequency value, $\mathbf{\hat{d}}$ is the predicted data for the given frequency component (including the real and imaginary parts), $\mathbf{v}_{l}$ is input to the ${l+1}^{th}$ FNO layer ($l=0,...,7$) or the GNO ($l=8$), $\mathbf{v}_{gno}$ is the GNO output, $\mathcal{P}$ is a point-wise operator used to lift the dimension, $\mathcal{Q}$ is used to project the output to the desired dimension, $\mathcal{W}_{l}$ acts as a residual connection, $\kappa_{l}$ is a parametric kernel function, $\mathcal{F}$ and $\mathcal{F}^{-1}$ denote the Fourier and inverse Fourier transforms, respectively, and $\sigma$ is a nonlinear activation function, which we set as GELU (Hendrycks & Gimpel, 2016). The GNO uses a kernel function $\kappa$ parameterized as a three-layer neural network, which takes positional information and the output of the last FNO layer as input. $B_{r}$ denotes the predefined neighborhood of the queried position $x$, defined as all points below the corresponding surface location for computational efficiency. We use FNOs to speed up the integral calculation, and the U-shape allows for a deeper model and skip connections (Li et al., 2020c; Rahman et al., 2022; Ronneberger et al., 2015). The GNO can improve predictions at queried points and provide greater data efficiency (Zou et al., 2024). Figure 1 illustrates the model architecture and Table S1 provides the details in each neural operator layer.
@@
where $\mathbf{\hat{d}}$ is the modeled (predicted) waveforms queried at receivers through a sampling operator $\mathcal{S}$ and $\mathbf{d}$ is the observed data . The parameters of interest $\mathbf{m}$ are updated iteratively using a gradient-based method. Using the adjoint-state method, the gradients of the objective with respect to parameters of interest can be estimated as the cross-correlation between the forward wavefield $\mathbf{u}$ and the adjoint wavefield $\mathbf{u}^{*}$ at zero time lag, weighted by an analytically derived term (Louboutin et al., 2019):
@@
The adjoint-state method requires the initial model to be sufficiently accurate, with a travel-time error of less than half the period (Beydoun & Tarantola, 1988; Virieux & Operto, 2009; Alkhalifah, 2016). Otherwise, the cycle-skipping phenomenon may arise, leading to convergence to a local minimum (Gauthier et al., 1986; Pladys et al., 2021).
With neural operators (or other machine learning methods developed on the PyTorch platform), the gradients of the objective with respect to any parameters in the computational graph are automatically computed through back-propagation (Rumelhart et al., 1986). This is termed the reverse-mode AD (Baydin et al., 2018; Elliott, 2018). Provided that all the functions, expressions, and control flow structures used in neural operators are differentiable and compatible with AD, the adjoint simulation can be replaced by back-propagation (LeCun et al., 1988). It has been shown that the adjoint-state method and reverse-mode AD are mathematically equivalent (Richardson, 2018; Wang et al., 2021b; Zhu et al., 2021). However, the HNO-AD method eliminates the need to manually derive the gradient in Equation 7 on a case-by-case basis. Moreover, it can take advantage of state-of-the-art optimization techniques built into PyTorch — such as mini-batching and Adam — which help improve the optimization dynamics of FWI and mitigate cycle-skipping. Adam, in particular, a first-order gradient-based optimization algorithm that dynamically rescales the gradients for each parameter based on its past gradients and their squares (Kingma & Ba, 2014), has been shown to outperform other optimizers such as SGD and limited-memory Broyden-Fletcher-Goldfarb-Shanno (L-BFGS) (Bernal-Romero & Iturrarán-Viveros, 2021; Richardson, 2018; Sun et al., 2020).
@@ 2.2 Data-Driven Training
In this study, we train a single 2D HNO for linear nodal arrays across the northern LA basins. The model is trained in a supervised learning manner using synthetic data generated from a spectral element method (SEM) using the software SALVUS (Afanasiev et al., 2019). The computational domain is set to 808080 km (horizontal) $\times$ 404040 km (vertical) on a 256×128256$\times$ 128256 × 128 mesh, which accommodates the longest seismic line, SB1, in the area. We simulate 404040-s-duration wavefields with a time step of 0.0010.0010.001 s, which are excited by a vertical force and Ricker wavelet time function with a central frequency of 0.30.30.3 Hz. For ambient noise data applications, the source is randomly placed along the free surface from a uniform distribution of horizontal position. The $V_{P}$ and $V_{S}$ models are generated from random fields sampled around a background 1D model, which is averaged along the SB1 line from a reference model named CVM-S4.26 (Lee et al., 2014). The parameters for the random fields are provided in Table S2. The density models are derived from $V_{P}$ using the empirical relation by Brocher (2005), which are input to SALVUS but not to the HNO. The time-domain solutions from the SEM solver are Fourier-transformed and filtered to the frequency band of interest (0.10.10.1 to 0.50.50.5 Hz) for use with the HNO. We train an HNO using 270002700027000 simulations and validate the model performance using 300030003000 simulations (where a simulation means a single source and a particular velocity model). We define the loss function as a combination of relative $\ell_{1}$-norm and $\ell_{2}$-norm of the data residual:
@@
where $\mathbf{d}$ and $\mathbf{\hat{d}}$ denote the true and predicted data, respectively. We use an Adam optimizer with a learning rate of 0.0010.0010.001 and a scheduler that decays the learning rate by half every 303030 epochs. The batch size is 256256256. Figure S1 shows the loss curves. The full training process took 222 days with 888 NVIDIA RTX A6000 GPUs. Note that the training is a one-time effort and no further training is required for different seismic lines or the inversion stage, granted the data have a similar scale. Table 1 compares the time consumed by SALVUS and the trained HNO in forward modeling and FWI under the same configuration. In performing FWI for this 2D case, the HNO with AD is approximately two orders of magnitude faster than the SEM with the adjoint method. The computational advantage will only be greater in 3D (Zou et al., 2024).
Table 1: Comparison of time consumed by SALVUS and HNO in forward modeling and FWI. The experiments are conducted on a community model called CVM-S4.26 using a single NVIDIA RTX A6000 GPU. The forward modeling time is measured for a single event with one source and 234234234 receivers on the free surface. The FWI time is measured for one tomographic iteration for the same event configuration.
@@
2.3 Synthetic Tests
To test generalization, we first evaluate the HNO performance with the synthetic data generated from a community velocity model named CVM-S4.26 (Lee et al., 2014). We extract a profile of the 3D CVM-S4.26 along the longest seismic line, SB1, in the BASIN survey. This model is never explicitly seen by the HNO. Figure 2 shows the HNO-predicted results for a sample shot excited by a vertical force and Ricker wavelet as in training, with the SALVUS-simulated waveforms serving as the ground truth. The direct output of the HNO is in the frequency domain (0.10.10.1 to 0.50.50.5 Hz) and is inverse Fourier-transformed to the time domain for visualization purposes. The forward prediction by the HNO is not perfect with a cross-correlation coefficient of 0.98, but we will show that it is sufficiently accurate in the upcoming FWI experiment.

### v4

@@ 1 Introduction
Seismic tomography for sedimentary basins is important for assessing earthquake hazards considering the fact that the basins can trap and amplify seismic waves, leading to stronger and longer ground shaking. Fine-scale subsurface structures are typically determined by active source surveys, which often operate at high frequencies but are expensive and limited by environmental impact issues. Alternatively, seismic data from passive sources, such as ambient noise and earthquakes, can be more readily acquired and can also compensate for the absence of low-frequency information in active source data. Many have shown in seismic interferometry that the cross-correlation between diffuse waveforms recorded at two stations approximates the elastodynamic Green’s function (wapenaar2004retrieving; wapenaar2010tutorial). The so-called empirical Green’s function (EGF) created from ambient noise is more commonly used by ray-theory-based tomography methods (guo2015high; lin2008surface; yao2010heterogeneity; zheng2011crust), which do not account for the finite-frequency effects in heterogeneous media. Full waveform inversion (FWI), which accounts for the full physics of wave propagation, can reveal more accurate and detailed subsurface structures (liu20173; liu2023ambient; maguire2022magma; sager2020global; wang2021adjoint; zhang2021rayleigh).
@@
Despite its significant advantages, FWI can be time-consuming and memory-intensive for large problems, which hinders its broader applicability. Moreover, the adjoint-state method (fichtner2006adjoint1; liu2006finite; plessix2006review; tromp2005seismic), which is conventionally used for FWI, requires the initial model to be sufficiently close to the true Earth to avoid being trapped in local minima (gauthier1986two; fichtner2006adjoint2; panning2009seismic; virieux2009overview). In ambient noise tomography (ANT), the adjoint-state method typically relies on a model derived from surface wave dispersion analysis and ray-based approximations as a starting point (chen2014low; liu2023ambient; maguire2022magma; zhang2018linear; zhang2020wave). To tackle these challenges, recent studies have shown that neural operators (li2020fourier; li2020neural; azizzadenesheli2024neural) can solve wave equations orders of magnitude faster than traditional numerical solvers (yang2021seismic; yang2023rapid; zou2024deep). In deep learning, an analogue to the adjoint-state method is known as the reverse-mode automatic differentiation (AD) (baydin2018automatic; rumelhart1986learning). While these two methods have been shown to be mathematically equivalent (richardson2018seismic; wang2021elastic; Zhu), AD, which underpins the modern deep learning platforms such as PyTorch (pytorch) and TensorFlow (TensorFlow), remains aligned with advances in state-of-the-art optimization techniques. For example, methods such as mini-batching (lecun2002efficient), stochastic gradient descent (SGD) (bottou2010large), and adaptive moment estimation (Adam) (Kingma) have been developed to better navigate non-convex loss landscapes and escape poor local minima. These methods have prevailed for over a decade in the machine learning community, while remaining relatively underexplored in the classic FWI world. Only in recent years have they begun to attract attention (bernal2021accelerating; mao2025automatic; richardson2018seismic; sun2020theory; thrastarson2022data; van2020accelerated). Although efforts have been made to implement finite difference solvers in PyTorch/TensorFlow to enable access to AD and various optimization strategies (richardson2018seismic; Zhu), this approach is memory prohibitive for realistic applications, because back-propagation requires storing wavefields at every time step in memory. We bypass this obstacle by learning a direct mapping through neural operators developed in PyTorch. In alignment with the rapid advancements in optimization techniques, neural operators have the potential to alleviate — though not fully resolve — the cycle-skipping problem in FWI, making it possible to start from a simpler velocity model (e.g., a 1D model).
@@
The primary advantage of neural operators in solving partial differential equations (PDEs) over other prevailing machine learning approaches, such as physics-informed neural networks (PINNs) (raissi2019physics), is that they are trained to learn the solution operator for an entire family of PDEs instead of a specific instance. This means that a neural operator has the potential to generalize to arbitrary PDE coefficients (e.g., velocity structures) if the training instances are ideally sampled — for example, from random fields (yang2021seismic; yang2023rapid; zou2024deep). The training is a one-time effort for the forward process and there is no further training for the neural operator in the inversion stage, where the velocity parameters are updated with gradients computed from AD. This is similar to traditional FWI but replaces the adjoint simulation with back-propagation (LeCun).
@@
In recent years, neural operators have been increasingly applied to seismic wave propagation modeling. They have been used for modeling acoustic (yang2021seismic), elastic (li2025feature; yang2023rapid), and viscoelastic (wei2022small) waves. Most applications are limited to 2D due to computational and memory constraints. However, lehmann20243d; lehmann2025multiple circumvented the 3D computational bottleneck by predicting only the surface ground motion, while zou2024deep modeled the full 3D wavefield by parameterizing the operator in the frequency domain, thereby eliminating the need to model the time dimension. Learning frequency-domain solutions offers memory efficiency for neural operators and has attracted growing interest (cheng2025seismic; huang2025learned; kong2025reducing; li2023solving; zhang2023learning). Among these, huang2025learned and cheng2025seismic proposed learning the residual (scattered) wavefield instead of the full wavefield to address the point-source singularity. Current research has primarily focused on synthetic data, while the application of neural operators to real seismic data remains largely unexplored.
@@
Our contributions are as follows. We demonstrate the first application of neural operators for ambient noise tomography on a real seismic dataset, specifically the Basin Amplification Seismic Investigation (BASIN) survey (BASIN2018; clayton2019exposing), which deployed linear nodal arrays in the northern Los Angeles (LA) basins. We provide a trained neural operator that is applicable without retraining to any linear array in the study area, or potentially other problems with a comparable spatial footprint. With the trained neural operator, FWI can be performed two orders of magnitude faster than the conventional adjoint method. Our tomography results are consistent with previous studies in the same region (ghose2023basin; li2023shear; villa2023three; Zou2024) without the strong dependence on priors. The proposed method can be scaled to 3D, provided that sufficient computing resources are available.
@@ 2 Methods
In this section, we first present the governing equations, as well as the methodology of using neural operators to solve the forward and inverse problems. Then, we share the training strategy.
@@ 2.1 Helmholtz Neural Operator with Automatic Differentiation
Based on the universal approximation theorem, neural operators can approximate arbitrary nonlinear continuous operators (hornik1989multilayer; Kovachki). The solution operator can be parameterized much more efficiently in memory when solving Equation 2 in the frequency domain (i.e., the Helmholtz equation) than in the time domain (huang2025learned; kong2025reducing; zou2024deep). This is because different frequency components can be solved individually and, thus, parallelized. In this study, we employ the Helmholtz Neural Operator (HNO) proposed by zou2024deep. The HNO computes the solution for each frequency component independently by defining the batch of data at the frequency level. In other words, a batch can contain multiple frequency components due to data shuffling. We use a single HNO to learn the entire frequency band of interest ($0.1$ to $0.5$ Hz), although kong2025reducing found that using multiple models to handle different frequencies slightly improves performance. Solutions in the time and frequency domains can be related via the (inverse) Fourier transform. We construct a U-shaped HNO comprising $L=8$ inner layers of Fourier Neural Operators (FNOs) (li2020fourier), followed by a Graph Neural Operator (GNO) (li2020neural), to predict the waveforms at the free surface (data, denoted by $\mathbf{\hat{d}}$):
@@
where $\mathbf{a}$ is the input including $V_{P}$, $V_{S}$, the source location, and a constant function indicating the frequency value, $\mathbf{\hat{d}}$ is the predicted data for the given frequency component (including the real and imaginary parts), $\mathbf{v}_{l}$ is input to the ${l+1}^{th}$ FNO layer (l=0,,7l=0,...,7) or the GNO ($l=8$), $\mathbf{v}_{gno}$ is the GNO output, $\mathcal{P}$ is a point-wise operator used to lift the dimension, $\mathcal{Q}$ is used to project the output to the desired dimension, $\mathcal{W}_{l}$ acts as a residual connection, $\kappa_{l}$ is a parametric kernel function, $\mathcal{F}$ and $\mathcal{F}^{-1}$ denote the Fourier and inverse Fourier transforms, respectively, and $\sigma$ is a nonlinear activation function, which we set as GELU (hendrycks2016gaussian). The GNO uses a kernel function $\kappa$ parameterized as a three-layer neural network, which takes positional information and the output of the last FNO layer as input. $B_{r}$ denotes the predefined neighborhood of the queried position $x$, defined as all points below the corresponding surface location for computational efficiency. We use FNOs to speed up the integral calculation, and the U-shape allows for a deeper model and skip connections (li2020multipole; Rahman; Ronneberger). The GNO can improve predictions at queried points and provide greater data efficiency (zou2024deep). Figure 1 illustrates the model architecture and Table S1 provides the details in each neural operator layer.
@@
where $\mathbf{\hat{d}}$ is the modeled (predicted) waveforms queried at receivers through a sampling operator $\mathcal{S}$ and $\mathbf{d}$ is the observed data . The parameters of interest $\mathbf{m}$ are updated iteratively using a gradient-based method. Using the adjoint-state method, the gradients of the objective with respect to parameters of interest can be estimated as the cross-correlation between the forward wavefield $\mathbf{u}$ and the adjoint wavefield $\mathbf{u}^{*}$ at zero time lag, weighted by an analytically derived term (devito):
@@
The adjoint-state method requires the initial model to be sufficiently accurate, with a travel-time error of less than half the period (beydoun1988first; virieux2009overview; Tariq2016). Otherwise, the cycle-skipping phenomenon may arise, leading to convergence to a local minimum (gauthier1986two; pladys2021cycle).
With neural operators (or other machine learning methods developed on the PyTorch platform), the gradients of the objective with respect to any parameters in the computational graph are automatically computed through back-propagation (rumelhart1986learning). This is termed the reverse-mode AD (baydin2018automatic; elliott2018simple). Provided that all the functions, expressions, and control flow structures used in neural operators are differentiable and compatible with AD, the adjoint simulation can be replaced by back-propagation (LeCun). It has been shown that the adjoint-state method and reverse-mode AD are mathematically equivalent (richardson2018seismic; wang2021elastic; Zhu). However, the HNO-AD method eliminates the need to manually derive the gradient in Equation 7 on a case-by-case basis. Moreover, it can take advantage of state-of-the-art optimization techniques built into PyTorch — such as mini-batching and Adam — which help improve the optimization dynamics of FWI and mitigate cycle-skipping. Adam, in particular, a first-order gradient-based optimization algorithm that dynamically rescales the gradients for each parameter based on its past gradients and their squares (Kingma), has been shown to outperform other optimizers such as SGD and limited-memory Broyden-Fletcher-Goldfarb-Shanno (L-BFGS) (bernal2021accelerating; richardson2018seismic; sun2020theory).
@@ 2.2 Data-Driven Training
In this study, we train a single 2D HNO for linear nodal arrays across the northern LA basins. The model is trained in a supervised learning manner using synthetic data generated from a spectral element method (SEM) using the software SALVUS (Afanasiev). The computational domain is set to $80$ km (horizontal) $\times$ $40$ km (vertical) on a $256\times 128$ mesh, which accommodates the longest seismic line, SB1, in the area. We simulate $40$-s-duration wavefields with a time step of $0.001$ s, which are excited by a vertical force and Ricker wavelet time function with a central frequency of $0.3$ Hz. For ambient noise data applications, the source is randomly placed along the free surface from a uniform distribution of horizontal position. The $V_{P}$ and $V_{S}$ models are generated from random fields sampled around a background 1D model, which is averaged along the SB1 line from a reference model named CVM-S4.26 (lee2014full). The parameters for the random fields are provided in Table S2. The density models are derived from $V_{P}$ using the empirical relation by brocher2005empirical, which are input to SALVUS but not to the HNO. The time-domain solutions from the SEM solver are Fourier-transformed and filtered to the frequency band of interest ($0.1$ to $0.5$ Hz) for use with the HNO. We train an HNO using $27000$ simulations and validate the model performance using $3000$ simulations (where a simulation means a single source and a particular velocity model). We define the loss function as a combination of relative $\ell_{1}$-norm and $\ell_{2}$-norm of the data residual:
@@
where $\mathbf{d}$ and $\mathbf{\hat{d}}$ denote the true and predicted data, respectively. We use an Adam optimizer with a learning rate of $0.001$ and a scheduler that decays the learning rate by half every $30$ epochs. The batch size is $256$. Figure S1 shows the loss curves. The full training process took $2$ days with $8$ NVIDIA RTX A6000 GPUs. Note that the training is a one-time effort and no further training is required for different seismic lines or the inversion stage, granted the data have a similar scale. Table 1 compares the time consumed by SALVUS and the trained HNO in forward modeling and FWI under the same configuration. In performing FWI for this 2D case, the HNO with AD is approximately two orders of magnitude faster than the SEM with the adjoint method. The computational advantage will only be greater in 3D (zou2024deep).
Table 1: Comparison of time consumed by SALVUS and HNO in forward modeling and FWI. The experiments are conducted on a community model called CVM-S4.26 using a single NVIDIA RTX A6000 GPU. The forward modeling time is measured for a single event with one source and $234$ receivers on the free surface. The FWI time is measured for one tomographic iteration for the same event configuration.
@@
3 Synthetic Tests
To test generalization, we first evaluate the HNO performance with the synthetic data generated from a community velocity model named CVM-S4.26 (lee2014full). We extract a profile of the 3D CVM-S4.26 along the longest seismic line, SB1, in the BASIN survey. This model is never explicitly seen by the HNO. Figure 2 shows the HNO-predicted results for a sample shot excited by a vertical force and Ricker wavelet as in training, with the SALVUS-simulated waveforms serving as the ground truth. The direct output of the HNO is in the frequency domain ($0.1$ to $0.5$ Hz) and is inverse Fourier-transformed to the time domain for visualization purposes. The forward prediction by the HNO is not perfect with a cross-correlation coefficient of 0.98, but we will show that it is sufficiently accurate in the upcoming FWI experiment.